The sum of the complement and the supplement of an angle is 148 degrees, what is the angle?

Answer by lwsshak3(11628)

The sum of the complement and the supplement of an angle is 148 degrees, what is the angle?
***
let x=angle
90-x=complement of the angle
180-x=supplement of the angle
..
90-x+180-x=148
2x=270-148=122
x=61˚